### v3

# Invariance under Structure Translation as the Origin of Host Immune Capacity Conservation from Noether’s Theorem

Abstract
The capacity to resist pathogens is recognized as a fundamental property of the immune system, yet the capacity itself remains a phenomenological concept and lacks a defined physical basis. Its fundamental entity, definition, and quantification are thus unresolved. Here, we address these questions by introducing a theoretical framework based on Lagrangian analytical mechanics, which recasts immune recognition as a dynamical system in an immunological state space. Generalized coordinates are used to describe the conformational states of immune receptors, and their evolution is governed by Euler-Lagrange equations constructed from the antigen-receptor interaction. Central to our theory is the identification of a continuous symmetry: the action remains invariant under specific translations within the antigenic structure space or time. From this symmetry, Noether’s theorem dictates a conserved quantity, $I$. We propose that $I$ is the physical embodiment of host immunity, a quantifiable measure that integrates the system’s protective sensitivity (with dimensions of action) and response intensity (with dimensions of energy). Furthermore, this framework unifies key immunological phenomena as dynamical consequences of the same underlying conservation law, including vaccination, immune memory, tolerance, original antigenic sin, and T cell exhaustion. The consistency of this model with established clinical observations (e.g., conserved symptom profiles across distinct influenza strains) and published experimental data provides its initial validation. By transforming immune capacity from a phenomenological concept into a quantifiable physical entity defined by a conservation law, this work establishes a foundational framework for a unified, predictive immunology.
Key Words: Theoretical Immunology; Systems Immunology; Theoretical Biophysics; Lagrangian Mechanics; Noether’s Theorem; Symmetry and Conservation Laws; Immune Capacity; Immune Repertoire
Author’s Note: This manuscript presents a theoretical contribution and a testable experimental prediction. The theoretical prediction find preliminary support in clinical observations, but remain phenomenological and lack the precise, quantitative criterion. The experimental work on model organisms in Result Section is currently in progress and will be reported in a future version. The list of authors is subject to change based on the experimental work. Comments and feedback on the theoretical framework are welcomed.
HighLights:
1. A new analytical mechanics formalism for immunology. Develops a Lagrangian formulation within an immunological state space, translating immune recognition and response into a principled dynamical system.
2. Immunity as a Conserved Quantity from Symmetry. Identifies translational invariance in antigenic structure and time as fundamental symmetries, from which Noether’s theorem yields a conserved physical quantity, $\mathbf{I}$, that defines immune capacity.
3. Decouples immune capacity into a generalized momentum (encoding sensitivity) and a generalized force power (encoding intensity), both derived from first principles.
4. Unification of Disparate Immune Phenomena. Provides a single dynamical explanation for vaccination, memory, tolerance, exhaustion, and original antigenic sin as distinct manifestations of the same conserved quantity $\mathbf{I}$ under different boundary conditions.
5. Validation Through Clinical and Experimental Consistency. The model naturally explains well‑established but previously disconnected observations, such as conserved symptom profiles across antigenically distinct strains, thereby bridging theoretical immunology with clinical phenomenology.
1 Introduction
The immune system constitutes a remarkably sophisticated defense network, capable of generating highly specific and potent responses against an immense diversity of pathogens. This protective capacity, broadly termed immune capacity, underpins organismal survival and is the cornerstone of vaccination strategies. While significant progress has been made in understanding molecular mechanisms like antibody-antigen binding kinetics, T-cell receptor signaling and the cellular dynamics of clonal selection, a predictive, first-principles theoretical framework remains a fundamental challenge. Current models largely operate phenomenologically, correlating specific molecular features(e.g., epitope affinity, antigen dose) with limited outcomes but failing to systematically predict the emergent, system-level immune capacity, which is an integrated metric of protection that encompasses both the sensitivity of recognizing antigens and the intensity of the elicited response. Critically, interventions aimed at ”enhancing immunity” are often described in reductionist terms, such as stimulating the proliferation of specific cell types. A unified theory is needed to describe how such molecular-scale perturbations systematically reconfigure the entire system to elevate the global protective capacity, and a first-principles framework that defines and quantifies this emergent capacity is therefore essential for advancing predictive immunology and rational intervention design.
The immune system’s protective function is an emergent property, arising from the collective interactions of innumerable cells and molecules. To describe such a system, the key variables cannot be the positions of individual atoms but must be coarse-grained, emergent order parametersthat capture the macroscopic state, much like ’temperature’ emerges from molecular motions but serves as a powerful coordinate in thermodynamics. Analytical mechanics, particularly the Lagrangian formalism, is uniquely suited for this task. Its power lies in describing complex systems using generalized coordinates, they are abstract parameters defined solely by the system’s degrees of freedom, independent of their microscopic physical nature. The resulting equations of motion are derived systematically from variational principles, providing a rigorous and unified dynamical description. This framework shifts the focus from mechanistic details of individual components to action principles and invariants, allowing it to capture the emergent dynamics that define immunological function. We therefore posit that the immune system represents a high-dimensional dynamical system amenable to Lagrangian description. Translating immune recognition into this framework offers the potential to uncover fundamental organizing principles beyond phenomenological models. In this work, we implement this translation to reveal a conserved physical quantity that underlies what is phenomenologically described as ”immune capacity”.
In this work, we establish a novel theoretical framework for adaptive immunity based on Lagrangian mechanics and symmetry principles. We begin by defining an immunological state space with generalized coordinates and constructing a Lagrangian, from which the system dynamics are derived via the Euler-Lagrange equations. Crucially, we identify a continuous symmetry, under which the action remains invariant: translational invariance in antigenic structure space. Through Noether’s theorem, this symmetry yields a conserved quantity, I. We demonstrate that I physically embodies what is phenomenologically termed “immune capacity,” providing it with a rigorous, first-principles definition as a conserved property of the dynamical system. This framework naturally decouples immune capacity into two components: a generalized momentum (encoding the sensitivity of recognition) and the power of the generalized force (encoding the intensity of the response). Importantly, the model offers a unified explanation for diverse immunological phenomena, including vaccination, memory, tolerance, exhaustion, and original antigenic sin as distinct dynamical consequences of the same underlying conservation law. The consistency of this theoretical picture with established clinical observations and published experimental data is discussed as a key validation. By transforming immune capacity from a phenomenological metaphor into a quantifiable conserved entity, this work provides a foundational framework for predictive immunology and rational intervention design.
2 Analytical Mechanics of Immune Recognition
2.1 Multi-scale Potential Energy Surface Theory
The immune response is initiated at the molecular level by the specific binding between an antigenic epitope and a cognate receptor (e.g., Toll Like receptor, B-cell or T-cell receptor). This physicochemical process is conceptualized within a potential energy landscape(Fig. 1a). In the unbound state, the receptor resides in a meta-stable energy minimum. Successful binding requires the system to overcome an activation energy barrier, a process driven by thermal fluctuations or cellular motions, leading to a conformational change and a transition to a more stable, lower-energy bound state (Fig. 1b). The height of this barrier governs the binding probability, whereas the energy difference ($\Delta E$) between the initial and final states is determined jointly by the intrinsic structural potential of the antigen and the receptor, as well as the coupling potential released upon their interaction as the system relaxes into a lower-energy state. This $\Delta E$ represents the driving force that propels downstream biochemical cascades, ultimately culminating in cellular activation, proliferation, and differentiation.
This molecular recognition event extends to the cellular population level. The adaptive immune system maintains a vast repertoire of receptors through stochastic genetic recombination. Consequently, the response to a specific antigen is not mediated by a single receptor type but by an ensemble of clones with a continuous distribution of binding affinities. In parallel, the innate immune system uniformly modifies the intensity of antigen signals in a non-specific manner. This results in a corresponding distribution of activation energy barriers across the responsive cell population, rather than a single value (Fig. 1c). The statistical properties of this distribution, including its mean, variance, and shape, collectively determine the population’s overall response intensity and sensitivity.
At the macroscopic scale, the immune system must be prepared to respond to a near infinite universe of potential antigens. It employs a three-pronged strategy: deploying a consistent innate immune response to handle low-level stimuli and mitigate the intensity of stronger challenges, generating pattern-recognition receptors (e.g., Toll Like Receptors, TLRs) with pre-defined, low-energy barriers for conserved pathogenic structures, while also maintaining an immense diversity of somatically rearranged receptors (e.g., T Cell Receptors, TCRs, and B Cell Receptors, BCRs) to cover unforeseen threats. This creates a high-dimensional recognition space where each potential antigen maps to a specific region, activating a unique subset of the total repertoire (Fig. 1d). Thus, the interaction between the antigenic universe and the immune system is no longer described by individual energy curves but emerges as a high-dimensional Potential Energy Surface (PES), which encodes the system’s entire landscape of reactivity.
To account for the open, dissipative nature of the biological system, where energy and matter are constantly exchanged with the environment (e.g., through cell death, proliferation, metabolic input), we introduce generalized forces ($Q$). The magnitude of $Q$ is regulated by factors such as the intensity of the endocrine system and is determined by the current state of the immune system. In the early stages of the immune response, $Q$ acts in an activating direction, driving the immune response toward greater intensity. In later stages QQbecomes inhibitory and dissipative, promoting the return of the immune response to homeostasis. The generalized forces can be regarded as intrinsic properties of the system’s instantaneous state.
The timescales of antigen recognition (on the order of milliseconds) are vastly shorter than those of systemic potential energy surface (PES) remodeling, which occurs over days. Consequently, during a single antigen recognition event, the macroscopic PES ($V$) can be considered approximately unchanged. For an acute immune response, the PES can thus be treated as static, and the system may be regarded as conservative. Moreover, when compared to extended processes such as aging, the body’s homeostatic condition remains relatively stable over the course of a single immune response cycle. For an individual of given age and health status, the generalized generalized force will take the same value provided the immune system gives identical signals ($Q$). In contrast, across longer-term immune processes, both the PES ($V$) and the generalized force $Q$ evolve through mechanisms such as aging, immune memory, and tolerance, thereby fundamentally reshaping the system’s dynamic behavior.
2.2 Definition of the Immune State Space and Generalized Coordinates
To mathematically capture the dynamics of this complex system, we employ the formalism of analytical mechanics. We define the system’s state using three orthogonal degrees of freedom as generalized coordinates, these coordinates are emergent order parametersthat capture the macroscopic state of the system, analogous to thermodynamic variables, and the generalized momenta are given by the partial derivatives of the Lagrangian with respect to the velocities, and the generalized masses, which phenomenologically quantify the system’s inertial resistance to changes in its state, are computationally determined by the partial derivatives of the momenta with respect to the velocities:
$p_{i}=\frac{\partial L}{\partial\dot{q_{i}}},m_{i}=\frac{\partial p_{i}}{\partial\dot{q_{i}}}$ (1)
$q_{1}$
: Immunological Repertoire Configuration. In single protein-antigen interaction, it represent the structure of the antigen-binding protein; In system level, it represent the distribution of total antigen-recognition states accessible to the system. The distribution is often measured as receptor repertoire diversity by Shannon entropy.
$q_{2}$
: Effective Antigenic Structure. A parameter that integrates key physicochemical properties of an antigenic epitope (e.g., charge distribution, hydrophobicity, steric topography) into a one-dimensional metric of its overall characteristics. It emerges from the complex intermolecular landscape, analogous to how temperature emerges from molecular motions.
$q_{3}$
:Cooperativity of Molecular Deformation, a parameter describing the degree of concerted conformational change undergone by the antigen-receptor complex during binding. It macroscopically captures the effects of induced fit and conformational selection.
$\dot{q_{1}}$
: The rate of change of the receptor repertoire configuration. encompassing processes such as clonal expansion, contraction, and the generation of new receptors, which dynamically alters the recognition landscape.
$\dot{q_{2}}$
: The rate of change of the antigenic structure. This velocity represents the antigenic drift or shift of a pathogen, quantifying the dynamical evolution of the antigen challenge. For an acute immune response to a static pathogen, $\dot{q_{2}}=0$. On evolutionary timescales, it quantifies antigenic drift.
$\dot{q_{3}}$
: Rate of Conformational Dynamics. The rate of structural adaptation during the antigen-receptor interaction, directly linked to the kinetics of molecular binding.
$m_{1}$
: The effevtive mass of immune system, represents the system’s inertial resistance to state change and is biologically correlated with the metabolic capacity of lymphoid tissues, in system level, it is also correlated with the total lymphocyte count and the structural integrity of the immune network.
$m_{2}$
: Effective Mass of Antigenic Structure, or the volutionary Inertia of the antigen, represents the inertial resistance of an antigenic structure to undergoing functionally viable mutations. It quantifies the evolutionary plasticity of a pathogen epitope, defining the difficulty of its spontaneous change.
$m_{3}$
: The Effective Mass of antigen receptor flexibility describes the inertial resistance to a conformational transition from a protein’s native state to a distinct functional or dysfunctional state. It reflects the intrinsic flexibility or rigidity of the interacting molecules.
Since cell apoptosis causes the immune repertoire to move spontaneously in the negative direction, whereas the entry of external antigens contributes a positive velocity, we adopt this convention to assign the signs for the generalized coordinates $q_{1}$ and $q_{2}$. A positive $\dot{q}_{1}$ corresponds to an increase in the proportion of the immune repertoire capable of recognizing the antigen under antigenic stimulation, while a negative $\dot{q}_{1}$ reflects the contraction due to homeostatic apoptosis; Similarly, a positive $\dot{q}_{2}$ indicates the increase of antigenic exposure, and a negative $\dot{q}_{2}$ represents antigen clearance or decay.
The generalized mass $m_{1}$ captures the inertia of the immune system as a whole, quantifying the resistance of its state to change under external antigenic stimuli. To bridge cellular level properties with system level dynamics, we express $m_{1}$ as a composite of three key factors: the total lymphocyte count $N$, the single cell equivalent inertia $\mu$, and the structural integrity of the immune network $\eta$. Here, $N$ quantifies the sheer size of the immune repertoire; $\mu$ reflects the intrinsic activation threshold and signaling efficiency of an individual lymphocyte, representing how readily a single cell alters its functional state; and $\eta$ measures the coherence of inter-cellular communication, lymphoid tissue organization, and overall coordination within the immune network. A value of small $\eta$ indicates various degrees of structural deterioration or functional decoupling. These three components combine multiplicatively and inversely to give:
$m_{1}=\frac{N\mu}{\eta}$ (2)
This formulation encapsulates the idea that a larger activated lymphocyte population $N$ increases the system’s inertia, whereas a higher network integrity $\eta$ reduces the effective inertia by enabling more coordinated, system wide responses. Consequently, the generalized mass $m_{1}$ not only depends on the number of cells but also on how well they are organized to act collectively, thereby providing a unified metric for the immune system’s resistance to perturbation at the organism level.
This set of coordinates and velocities spans the state space of the immune system, where the PES $V(q)$, a function of these coordinates, dictates the system’s dynamics, enabling a complete analytical mechanics description of immune recognition and response. The kinetic energy $T$ represents the system’s activity associated with changes in these emergent coordinates. It is formulated as:
$T=\frac{1}{2}\sum_{i=1}^{3}m_{i}(\frac{\mathrm{d}q_{i}}{\mathrm{d}t})^{2}$ (3)
The coupling potential term $V_{coupling}$ is jointly determined by the immune receptor repertoire ($q_{1}$) and the specific antigenic structure ($q_{2}$). Different antigenic structures engage distinct receptor clonotypes, thereby altering the total number and diversity of participating immune cells and shaping a distinct potential energy surface (PES). Because immune recognition is often pattern-directed rather than conformation-specific, multiple antigen-immune system combinations can yield the same potential energy distribution, for example, the cGAS protein primarily recognizes the phosphate-deoxyribose backbone of DNA and is largely insensitive to the nucleotide sequence itself, so that a single base change (e.g., from Adenine to Cytosine) usually does not affect its activation strength provided other parameters such as antigen amount and length remain unchanged. This functional relationship can be expressed as $V_{coupling}=f(G(q_{1},q_{2}))$, where Grepresents the effective immunogenicity/antigenicity function that quantifies the strength of the immune response elicited by the antigen-receptor interaction. The condition $G(q_{1},q_{2})=\text{constant}$ defines an “immunological equipotential surface” on which all antigen-receptor pairs exhibit the same immune-stimulatory efficacy. Its partial derivatives define the sensitive parameters:
$K_{1}=\frac{\partial G}{\partial q_{1}},K_{2}=\frac{\partial G}{\partial q_{2}}$ (4)
where $K_{1}$ measures the sensitivity to receptor structure variations, and $K_{2}$ measures the sensitivity of immunogenicity to antigenic structure variations.
2.3 The Lagrangian Equation of Motion
The evolution of the system within this state space is governed by the principles of analytical mechanics. The kinetic energy, $T=T(\dot{q_{i}})$ captures the dynamics associated with changes in the generalized coordinates. while the potential energy, $V=V(q_{i})$ represents the macroscopic within its structure
These forces phenomenologically represent biological processes such as clonal expansion (a driving force) and cell death (a dissipative force). This leads to the Lagrangian expression and equation of motion:
$L=\Sigma T(\dot{q_{i}})-\Sigma V(q_{i})-V_{\text{couple}}(G(q_{1},q_{2}))$ (5)
$\frac{\mathrm{d}}{\mathrm{d}t}(\frac{\partial L}{\partial\dot{q_{i}}})-\frac{\partial L}{\partial q_{i}}=0$ (6)
The system’s trajectory will spontaneously evolve according to the constraints embedded in this Lagrangian.
Considering the entire immune process and include $Q(q_{1})$, the Lagrangian then become:
$\frac{\mathrm{d}}{\mathrm{d}t}(\frac{\partial L}{\partial\dot{q_{i}}})-\frac{\partial L}{\partial q_{i}}=Q(q_{1})$ (7)
2.4 Interpreting Immune Phenomena
The established dynamical equation provides a unified physical interpretation for core immunological phenomena:
Affinity
describes single-bond strength, corresponds to a deep and narrow potential well on the PES.
Avidity
results from multivalent interactions, arises from the spatial summation of multiple wells, which collectively lowers the effective activation barrier.
Immune Tolerance
is physically implemented as an extremely high activation barrier in the PES region corresponding to self-antigens. This makes the action for self-reactive pathways approach infinity, fundamentally preventing activation.
Immune Memory
is a beneficial deformation of the PES following infection or vaccination, characterized by lowered barriers and deepened wells for a specific pathogen, leading to a minimized action for a secondary response in accordance with the principle of least action.
T-cell Exhaustion
can be modeled as a pathological reshaping of the PES under chronic stimulation, where functional wells become shallow or even replaced by repulsive barriers, reflecting a state where dissipation overwhelms driving forces.
Immunodominance
stems from the innate inhomogeneity of the PES, where immunodominant epitopes reside in regions with naturally lower barriers and deeper wells, making their pathways the ones of least action.
Original Antigenic Sin
exemplifies path dependence in the system’s dynamics. The memory wells carved by prior infections create deep minima in the action landscape. When encountering a similar antigen, the system’s evolution is constrained to paths proximal to this historical, low-action trajectory, thereby limiting the exploration of and response to novel epitopes.
These examples demonstrate that diverse immune behaviors, from molecular interactions to system-level responses, can be coherently explained by a unified PES topology and its governing dynamics. This confirms the framework’s powerful explanatory and unifying capacity.
2.5 Potential Degeneracy of Immune Recognition
Grounded in the symmetry principles of analytical mechanics, our theoretical framework leads to two fundamental and testable predictions regarding the degeneracy of immune recognition. First, spatial degeneracy predicts that if two distinct antigenic epitopes share identical coupling interactions with the immune system, resulting in the same effective barrier distributions and $\Delta E$, despite differing structural parameters $q_{1}$, $q_{2}$, they will elicit statistically indistinguishable immune responses. Second, temporal invariance predicts that over short timescales where core immune parameters($Q$) remain stable (neglecting longer-term cycles or drift), the system’s response to an identical challenge will be statistically consistent.
These theoretical predictions find preliminary support in common clinical observations. For instance, the diagnosis of respiratory tract infections often presents a dilemma: etiologically distinct pathogens (e.g., influenza virus vs. respiratory syncytial virus, or rhinovirus vs. seasonal coronavirus) can produce strikingly similar symptomatic profiles (e.g., fever, cough, malaise). Routine laboratory tests frequently only suggest a viral etiology without pinpointing the specific pathogen, hinting at a potential functional degeneracy at the system level.
This phenomenon is further corroborated by rigorous animal studies and omics analyses, which have consistently demonstrated that while two mice within the same experimental group may exhibit identical symptomatic manifestations, their generated B-cell and T-cell receptor (BCR/TCR) repertoires are never entirely identical. This observation provides strong evidence that immune response consistency is determined not by clonal identity at the single-sequence level, but by the statistical equivalence of the overall receptor distribution, which ultimately leads to indistinguishable functional outcomes.
However, these clinical observations remain phenomenological and lack the precise, quantitative criterion provided by our symmetry principle. Therefore, moving beyond these vague indications to rigorously test the specific hypotheses of spatial degeneracy and temporal invariance constitutes a critical validation step for our framework.
3 Experimental Framework and Testable Predictions
The model yields specific and testable predictions, to empirically test the spatial degeneracy prediction, a critical experiment would involve immunizing model organisms.
The experiment in this chapter are in progress.
3.1 Selection Criteria for Antigen Pairs
To probe the core predictions of functional degeneracy, we employe bacteriophage-derived virus-like particles (VLPs) as physiologically relevant antigens. Their stable, self-assembled spherical architecture permits precise structural perturbations via site-directed mutagenesis while preserving global integrity (e.g., size, stability, quaternary folding). This contrasts sharply with flexible peptide antigens, where even minor modifications risk uncontrolled conformational drift.
We engineer two key antigen pairs:
Degeneracy Test Pair (VLP-A/B): Differing by a single-point surface mutation, replacing the 38th amino acid Alanine with a Serine(A38S). This substitution introduces only a single hydroxyl group ,represents a minimal physicochemical perturbation while maintaining near-identical steric volume, polarity, and local hydrophilicity. This pair embodies the ”structurally similar but non-identical” antigens ideal for testing degeneracy.
Symmetry-Breaking Probe (VLP-C): Incorporating multiple cumulative mutations (F108L, A114K, S119D, L121M, Q128N). This design introduces significant, non-compensated alterations in charge distribution, hydrophobicity, and steric bulk, deliberately exceeding the threshold predicted for functional degeneracy.
3.2 Assessment of Immune Response Indistinguishability
Groups of C57BL/6 mice are immunized intraperitoneally (i.p.) with VLP-A, VLP-B, or VLP-C. To assess functional degeneracy, we measure immune responses across critical phases: First, the early innate activation (24h post-immunization): Activation states of dendritic cells and macrophages in draining lymph nodes are quantified to determine if initial antigen processing and innate immune triggering are indistinguishable between VLP-A and VLP-B, while potentially diverging for VLP-C. Second, Primary Adaptive Response (Day 7): The distribution and activation status of antigen-specific B and T cell populations, gene expressions with omics, alongside secreted antibody affinity profiles, are analyzed to test for statistical equivalence between VLP-A and VLP-B responses. Third, Secondary Response & PES Remodeling (Day 28 boost + Day 35 analysis): Following homologous boosting, memory responses are evaluated. Crucially, we assess whether the potential energy surface (PES) topology had remodeled between primary and secondary responses by comparing the kinetic pathways of B cell clonal expansion and epitope recognition breadth. Consistent responses to VLP-A/B across both primary and secondary challenges would confirm degeneracy, while divergent responses to VLP-C would demonstrate symmetry breaking under significant structural perturbation. the experiment with VLP-A are be repeated separately in 28-day-old and 35-day-old mice to examine the consistency of the aforementioned outcomes. This aims to validate whether a short temporal difference (7-day age gap) leads to temporal equivalence in immune state and whether immune responses remain consistent despite minimal variation in age.
This design rigorously tests the prediction and defines the limits of the symmetry principle.
3.3 Results
Experimental results and figures will be discussed here.
4 From Degeneracy to Symmetry: The Conservation of Immune Response
We posit that the observed consistency of immunity is not coincidental but must be governed by a fundamental physical principle. We now demonstrate how this symmetry mandates the existence of conserved quantities via Noether’s theorem, which we identify as the physical embodiment of Immune Capacity.
4.1 Spatial Continuous Symmetry
Our starting point is the Lagrangian formulation of the immunological dynamics. The state of the system is described by the generalized coordinates $(q_{1},q_{2},q_{3})$, the system’s dynamics are encoded in a Lagrangian $L=T(\dot{q_{i}})-V(q_{i})-V_{\text{couple}}(G(q_{1},q_{2}))$. During a single immune recognition event, the antigenic structure $q_{2}$ is fixed, and the immune state is held constant due to the conservative approximation; consequently, only the kinetic energy $T$ and the coupling potential $V_{\text{couple}}$ evolve over time. The actual path taken by the system between two times, and is the one that minimizes the action $S=\int L\,\mathrm{d}t$.
The symmetry can be explained by an invariance in the system’s dynamics. We postulate that the action $S$ remains invariant under a continuous transformation involving the coupling between antigenic epitope structure and the corresponding immune repertoire.
Consider a transformation that smoothly translates the system in the $q_{1}$ and $q_{2}$ coordinate by an infinitesimal amount $q_{1}\to q_{1}+\delta q_{1},q_{2}\to q_{2}+\delta q_{2}$, The symmetry principle is mathematically expressed by the requirement that the action remains invariant under this transformation:
$L^{\prime}=L(q_{1}+\delta q_{1},q_{2}+\delta q_{2},q_{3},\dot{q_{1}},\dot{q_{2}},\dot{q_{3}})$ (8)
The requirement of indistinguishability of the system under the transformation implies that the coupling between $q_{1}$ and $q_{2}$ remains unchanged, the value of $G$ is preserved. We therefore introduce such transformation where $\epsilon$ is an infinitesimal amount:
$G(q_{1},q_{2})=G(q_{1}+\delta q_{1},q_{2}+\delta q_{2})$ (9)
$\nabla G\cdot\delta\mathbf{q}=\frac{\partial G}{\partial q_{1}}\delta q_{1}+\frac{\partial G}{\partial q_{2}}\delta q_{2}=0$ (10)
$\delta q_{1}=\epsilon\frac{\partial G}{\partial q_{2}}=\epsilon K_{2},\delta q_{2}=-\epsilon\frac{\partial G}{\partial q_{1}}=-\epsilon K_{1}$ (11)
The variation of the action is
$\delta S=\int_{t_{1}}^{t_{2}}\delta L\mathrm{d}t=\int_{t_{1}}^{t_{2}}\left(\frac{\partial L}{\partial q_{1}}\delta q_{1}+\frac{\partial L}{\partial q_{2}}\delta q_{2}+\frac{\partial L}{\partial\dot{q}_{1}}\delta\dot{q}_{1}+\frac{\partial L}{\partial\dot{q}_{2}}\delta\dot{q}_{2}\right)\mathrm{d}t$ (12)
Assuming the transformation is a symmetry, $\delta S=0$, it follows that
$\delta S=\int_{t_{1}}^{t_{2}}\left(\frac{\partial L}{\partial q_{1}}\delta q_{1}+\frac{\partial L}{\partial q_{2}}\delta q_{2}+\frac{\partial L}{\partial\dot{q}_{1}}\cdot 0+\frac{\partial L}{\partial\dot{q}_{2}}\cdot 0\right)\mathrm{d}t=0$ (13)
$\delta S=\epsilon\int_{t_{1}}^{t_{2}}\left(\frac{\partial L}{\partial q_{1}}\frac{\partial G}{\partial q_{2}}-\frac{\partial L}{\partial q_{2}}\frac{\partial G}{\partial q_{1}}\right)\mathrm{d}t=0$ (14)
this condition leads to the conservation law:
$\frac{\partial L}{\partial q_{1}}\frac{\partial G}{\partial q_{2}}-\frac{\partial L}{\partial q_{2}}\frac{\partial G}{\partial q_{1}}=p_{1}K_{2}-p_{2}K_{1}=\text{Constant}$ (15)
$K_{1}K_{2}\left(\frac{p_{1}}{K_{1}}-\frac{p_{2}}{K_{2}}\right)=\text{Constant}$ (16)
where $p_{1}$ and $p_{2}$ are the generalized momenta. This conserved quantity arises from Noether’s theorem applied to the symmetry generated by the coupling function $G$. The equation implies that the quantity $I_{\mathrm{Spatial}}\equiv p_{1}K_{2}-p_{2}K_{1}$ is conserved. This is not an a priori assumption, but the simplest dynamical explanation for the observed phenomenological degeneracy.
After the short immune recognition progress, the generalized force $Q$ begins to affect the system, its introduction breaks the strict conservation of $I_{\mathrm{Spatial}}$.
4.2 Temporal Continuous Symmetry
When considering longer-term immune processes, the system transitions to an open system, wherein the total energy is governed by non-conservative generalized forces $Q$. The Hamiltonian canonical equations are given by:
$H=\sum_{i=1}^{n}p_{i}\dot{q}_{i}-L(q,\dot{q},t)$ (17)
$\dot{q}_{i}=\frac{\partial H}{\partial p_{i}}$ (18)
$\dot{p}_{i}=-\frac{\partial H}{\partial q_{i}}$ (19)
The time evolution of the Hamiltonian is expressed as:
$H=T+V+Q(q_{1})\dot{q}_{i}$ (20)
$\frac{\mathrm{d}H}{\mathrm{d}t}=Q(q_{1})\dot{q}_{i}$ (21)
Considering the energy change from time $t_{1}$ to $t_{2}$:
$\Delta H=\int_{t_{1}}^{t_{2}}\frac{\mathrm{d}H}{\mathrm{d}t}\ \mathrm{d}t=\int_{t_{1}}^{t_{2}}Q(q_{1})\dot{q}_{i}\ \mathrm{d}t$ (22)
This result implies that for systems starting from identical configurations and undergoing identical immune system changes($\Delta q$ and $\Delta\dot{q}$), the energy change $I_{\mathrm{Temporal}}\equiv\Delta E=\int_{t_{1}}^{t_{2}}Q(q_{1})\dot{q}_{i}\ \mathrm{d}t$ over identical time intervals $\Delta t$ remains invariant throughout the evolutionary process. This theoretical prediction aligns with our clinical observations, in which consistent immune responses were measured even in the presence of time-dependent external perturbations.
An alternative approach would involve expanding the system boundaries to incorporate the approximately conserved $Q(q_{1})$ parameters as internal degrees of freedom. However, this method would necessitate modeling additional complex subsystems like endocrine system, significantly increasing computational complexity without providing substantial physical insight. Therefore, we maintain the treatment of these parameters as external generalized forces, which adequately captures the essential dynamics while maintaining model tractability.
Through the application of analytical mechanics, we have derived a conserved quantity, $I_{\mathrm{Spatial}}=K_{2}p_{1}-K_{1}p_{2}$, and an invariant $I_{\mathrm{Temporal}}\equiv\int_{t_{1}}^{t_{2}}Q(q_{1})\dot{q}_{i}\ \mathrm{d}t$ directly from the symmetry inherent in the immune recognition process. This mathematical result provides a fundamental constraint governing the dynamics of the host-pathogen system. The existence of these conservation law is independent of specific pathological contexts; it is a general principle arising from the postulated structure of the interaction Lagrangian. The critical task that follows is to elucidate the profound biological implications of this constraint for immunology.
5 The Properties and Dynamic Responses of the Immune Capacity
The quantity derived in the preceding section embodies a fundamental property of the immune system. The system’s overall protective power is conserved and distributed across potential challenges. While acute responses to individual pathogens may vary, the global capacity remains constant for a given host state, bridging a fundamental physical principle with the core biological concept of host immunity. Therefore, we identify and named this quantity as the Immune Capacity $\mathbf{I}$:
$\mathbf{I}\equiv(I_{\mathrm{Spatial}},I_{\mathrm{Temporal}})$ (23)
$I_{\mathrm{Spatial}}=K_{2}p_{1}-K_{1}p_{2}$ (24)
$I_{\mathrm{Temporal}}=\int_{t_{1}}^{t_{2}}Q(q_{1})\dot{q}_{i}\ \mathrm{d}t$ (25)
The quantity $I_{\mathrm{Spatial}}$ defines the robustness of immune system to the varying antigen, and $I_{\mathrm{Temporal}}$ defines the intensity of an activated immune process. $\mathbf{I}$ integrates several static parameters that define the system’s inherent capacity, in the following sections, we first understand its fundamental properties in a steady state, then turn to the dynamic behavior of the immune system under the constraint of $\mathbf{I}$ conservation. This framework explains the wide spectrum of clinical responses to similar challenges.
5.1 The Dimensionality of Immune Capacity
To determine the dimensions of the Immune Capacity $\mathbf{I}$, we must first define the dimensions of the generalized coordinate $q_{1}$ and $q_{2}$. We posit that $q_{1}$ shows the diversity of immune repertoire and is dimensionless; $q_{2}$ is also an emergent, dimensionless order parameterthat integrates key physicochemical properties of the antigenic epitope, such as charge distribution, hydrophobicity patterns, and steric topography, into a single quantitative measure of ntigenic identity. Consequently, a translation in antigenic space, $q_{1}\to q_{1}+\delta q_{1}$, $q_{2}\to q_{2}+\delta q_{2}$, is a dimensionless operation.
The quantity $I_{\mathrm{Spatial}}$ therefore has the same fundamental dimension $\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-1}$ as angular momentum and action.
$I_{\mathrm{Spatial}}]=[p]=\frac{[E]\times[t]}{[q_{1}]}=\frac{[\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-2}\times[\mathrm{T}]}{1}=[\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-1}$ (26)
The dimension of the generalized force $Q$ is dictated by the principle of virtual work, $\delta W=Q\delta q_{1}$ must possess the dimensions of energy ($\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-2}$), Consequently, the dimensions of $Q$ are constrained by those of the corresponding virtual displacement $\delta q_{1}$, Specifically, if the generalized coordinate $q_{1}$ is dimensionless, the associated generalized force $Q$ must itself carry the dimensions of energy to ensure dimensional consistency of the virtual work expression. It therefore follows that $I_{\mathrm{Temporal}}$ possesses the dimensions of energy.
$I_{\mathrm{Temporal}}]=[\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-2}$ (27)
$\mathbf{I}$ quantifies the magnitude of the system’s persistent capacity. Much like angular momentum constitutes the charge arising from rotational symmetry and measures the persistent amount of rotation in a system, $I_{\mathrm{Spatial}}$ embodies a quantity stemming from a distinct symmetry inherent to the system, providing a stringent constraint on the system’s evolutionary pathways. While angular momentum is tied to the isotropy of space, $I_{\mathrm{Spatial}}$ is associated with a symmetry in antigenic epitope structure space.
While our phenomenological model treats $q_{2}$ as an emergent dimensionless parameter, its fundamental origin lies in the quantum mechanical electronic structure and nuclear coordinates of the antigen. At this most fundamental level, a translation in antigenic space would correspond to a collective rearrangement of atomic positions and electron cloud densities. However, this first-principles description is computationally intractable for biological systems. Our coarse-grained approach, which captures the emergent symmetry, is both necessary and sufficient for a systems-level immunological theory.
5.2 The Role of Innate Immunity within the Conservation
Within the established analytical mechanics framework of the immune response, characterized by the quantity $\mathbf{I}(I_{\mathrm{Spatial}},I_{\mathrm{Temporal}})$, innate immunity and other host systems are not represented by explicit dynamical variables. Instead, their influence is modeled as a modulation of key system parameters. Specifically, a robust innate immune response modulates the antigen stimuli $q_{2}$, the coupling function $G$ and generalized force $Q(q_{1})$. It first diminishes the effective antigenic load, thereby reducing the pathogenic impact that the adaptive immune system must counter, then it enhances the efficiency of antigen presentation and co-stimulatory signaling, which is quantified by an increase in the coupling function $G$. Concurrently, host metabolic and signaling activities underpin the generalized force $Q(q_{1})$, which represents the energetic cost required to maintain immune homeostasis and the degree of deviation from this steady state during disease. Thus, the dynamical evolution governed by the conservation law $\mathbf{I}(I_{\mathrm{Spatial}},I_{\mathrm{Temporal}})$ is fundamentally shaped by $G$, and $Q$, These parameters collectively establish the initial conditions and determine the operational trajectory of the immune response along the constraint defined by the conservation law.
5.3 Static Properties and Biological Interpretation of the Conserved Quantities
The quantity $I_{\mathrm{Spatial}}$ arises from the translational invariance of the generalized coordinates. It therefore constitutes a conservation law analogous to momentum conservation, quantifying the resistance to altering the immune system’s dynamical state during antigen interaction. An increase in the cell activation barrier $\mu$ or a decrease in immune-network integrity $\eta$ elevates this conserved quantity, reflecting greater difficulty in changing the system’s state. This manifests biologically as impaired immune activation and a sluggish response. Appropriate antigenic stimulation can reduce the magnitude of this conserved quantity, thereby lowering the activation threshold of the immune system, while excessive stimulation drives it too low. An excessively small value of $I_{\mathrm{Spatial}}$ indicates that the system’s state can be altered too readily, often correlating with hypersensitivity or autoimmune reactions. These dynamic adjustments are consistent with existing experimental observations. An optimal, intermediate magnitude representing a balanced level of resistance is therefore associated with appropriate immune responsiveness.
Similarly, $I_{\mathrm{Temporal}}$ originates from temporal translational invariance, leading to a consistency in energy variation. It characterizes the invariance of energy transfer during the immune system’s evolution under a given generalized force $Q$. When this quantity is too low, the open immune system receives an insufficient net energy influx, resulting in a weak response that may fail to clear antigens efficiently and can lead to chronic infection. If the value is too high, the immune reaction becomes excessively vigorous, potentially triggering pathological outcomes such as cytokine storms.
5.4 The Immune Dynamics under Impulse-Response Perturbation
The introduction of a significant pathogenic challenge imposes a positive impulse $\Delta p_{2}>0$ to the system. $I_{\mathrm{Spatial}}$ mandates an instantaneous compensatory adjustment such that $\Delta I_{\mathrm{Spatial}}=K_{1}\Delta p_{2}-K_{2}\Delta p_{1}=0$, leading to the fundamental constraint:
$K_{1}\Delta p_{2}=K_{2}\Delta p_{1}=\frac{K_{2}N\mu\Delta\dot{q}_{1}}{\eta}$ (28)
$\dot{q}_{1}=\frac{K_{1}\eta}{K_{2}N\mu}\Delta p_{2}$ (29)
Theoretically, this constraint predicts a positive generalized velocity $\dot{q}_{1}$. This prediction is consistent with published evidence showing that activated immune system during severe disease states has a higher expansion rate than that in milder diseases and healthy conditions.
The magnitude of the immune system’s response, quantified by the generalized velocity $\dot{q_{1}}$, is therefore determined by the magnitude of the pathogenic impulse $\Delta p_{2}$ and the system’s pre-existing total $I_{\mathrm{Spatial}}$. A large $\Delta p_{2}$ necessarily forces a substantial change in the immune state.
The host can achieve the requisite compensatory change $k\Delta p_{1}$ through several mechanistic strategies that modulate the parameters of the conservation equation. These include modulating the coupling efficiency $K_{1}$ through enhanced antigen presentation , altering the co-stimulation and inter-cell communication efficiency $\eta$, or directly reducing the antigenic load $p_{2}$ to diminish the effective $\Delta p_{2}$.
Based on the dynamical framework derived above, the change in the immune system’s state velocity $\dot{q}_{1}$, following an antigenic impulse $\Delta p_{2}$, is given by equation (29). Thus, the magnitude of $\dot{q}_{1}$ is directly proportional to the receptor marginal affinity K2, the antigenic impulse $\Delta p1$, and the network structural integrity $\eta$, while inversely proportional to the antigen marginal immunogenicity K1, the total lymphocyte number N, and the single-cell inertial parameter $\mu$. Among these factors, the parameters K1, K2, $\eta$, and $\mu$ typically exhibit relatively small inter-individual variation in a given population, whereas the total lymphocyte count Ncan vary substantially between individuals (e.g., young vs. elderly, healthy vs. immunocompromised). Consequently, a robust, healthy immune system with a large N experiences only a small change in its state velocity for a given pathogenic impulse; such a system can absorb the disturbance with minimal reconfiguration, maintaining homeostasis. In contrast, a weaker system with a smaller Nmust undergo a larger change in its velocity and potentially disruptive (i.e., a more drastic alteration of its repertoire composition or activation state) to satisfy the conservation law. This result predicts that the observed lineage-turnover rate or the speed of immune-repertoire reshaping should be inversely correlated with the overall immune competence, a prediction consistent with clinical observations that immunosenescence or immunodeficiency is often accompanied by exaggerated and dysregulated immune responses to infections.
Therefore a violent shift in the immune repertoire responding to a slight perturbation is not a sign of strength but rather an indicator of a system with low inertia and limited spare capacity, struggling to maintain homeostasis against a challenge.
Based on the antigenic impulse $p_{2}$, the system is driven from its homeostatic equilibrium. This deviation subsequently modulates the generalized force $Q$, which performs work on the immune system, thereby amplifying the positive displacement from equilibrium. The magnitude of this deviation at time $t$ is quantified by $I_{\mathrm{Temporal}}$. As antigen clearance progresses, the antigenic momentum $p_{2}$ declines. The reduction in $p_{2}$ induces a corresponding change of $Q$. Eventually, $Q$ may reverse direction, initiating a restorative phase that diminishes the system’s deviation and ultimately returns it to the baseline homeostatic state.
5.5 Integrating the Components as Total Immune Capacity
The total Immune Capacity $\mathbf{I}(I_{\mathrm{Spatial}},I_{\mathrm{Temporal}})$ emerges from the integration of its components, representing a holistic measure of the host’s protective potential. It is this integrated quantity but not any single parameter in isolation that most accurately corresponds to the clinical notion of Immune Capacity. A suitable value of $\mathbf{I}$ can result from various combinations of its constituents: a large total cell number $N$, a complete immune communication network $\eta$, an appropriate threshold for immune activation signals $\mu$, a highly coordinated multi-organ response (represented by $Q$), or a robust innate immune response for the preemptive reduction of antigenic load, which attenuates the impulse $\Delta p_{2}$ presented to the adaptive system under an identical challenge. This formalism quantifies the system-level sensitivity and amplitude of the immune state, indicating that while acute responses to specific pathogens may vary, the global protective capacity is maintained for a given host condition.
6 The Measurement of Immune Capacity
The previous section established the relationship between the conserved quantity for immune recognition, $I_{\mathrm{Spatial}}$, the invariant for response intensity, $I_{\mathrm{Temporal}}$, and key parameters ($k,N,\eta,\mu,\dot{q}_{1},Q$). Based on this theoretical framework, the subsequent critical task is the practical measurement or inference of these parameters and the Immune Capacity $\mathbf{I}$.
6.1 Determination of Immune Parameters in Model Organisms
The primary advantage of using model organisms lies in the ability to measure baseline parameters in a pristine state (devoid of external antigens) or under prolonged stability, and to observe systemic responses to standardized antigenic perturbations.
An experimental paradigm can be designed as follows: First, under homeostatic conditions (no antigenic challenge), $\dot{q}_{2}$ is zero, and the generalized force $Q$ is zero, indicating a balance between immunological energy dissipation and input. The conserved quantity simplifies to:
$I_{\mathrm{Spatial}}=k_{2}p_{1}=\frac{K_{2}N\mu\dot{q_{1}}}{\eta}$ (30)
Measuring the basal turnover rate of immune diversity $\dot{q}_{1}$, the existing total immune cell number $N$, the intrinsic activation threshold and signaling efficiency $\mu$ and the structural integrity of the immune network $\eta$, allows for an estimation of the fundamental Immune Capacity, $I_{\mathrm{Spatial}}$. For instance, parameters in a reference immune state (e.g., a healthy 42-day-old mouse) can be normalized ($N=1,\mu=1,\dot{q_{1}}=1,\eta=1$) for initial quantification.
Subsequently, applying a standardized antigenic challenge (e.g., VLP-A in this study), with a defined immunogenicity $G$, a standard $K_{1}=1,K_{2}=1$, and a impulse $p_{2}=1$ enables dynamic monitoring of the changes in $\dot{q}_{1}$, $\Delta p_{1}$, and the estimation of $I_{\mathrm{Temporal}}$ over time, thereby establishing a standard baseline of the systemic response.
This foundational approach allows for the assessment of disease models or therapeutic interventions. By comparing the altered immune dynamics under pathological or treated conditions against the baseline and the standard response curve, this model organism-based paradigm provides a robust basis for quantifying the efficacy of various interventions.
6.2 Assessment of Immune Capacity in Healthy Human Populations
When assessing Immune Capacity $\mathbf{I}$ in healthy human populations, the constant presence of environmental antigens must be considered. Continuous antigen exposure activates the immune system, resulting in non-zero values of the antigenic stimulation $q_{2}$ and $\dot{q}_{2}$ , which introduces a small yet persistent generalized momentum $p_{2}$. This in turn influences $p_{1}$ and $\dot{q}_{1}$, ultimately leading to a sustained displacement in $q_{1}$. Under conditions of persistent antigenic pressure, the system eventually reaches a dynamic equilibrium where the net generalized force returns to zero ($Q=0$), signifying a balance between energy dissipation within the immune system and energy input from antigenic stimulation. Consequently, the immune system resides in a state that is slightly yet stably displaced from a theoretical, antigen-free homeostasis.
Given the practical and ethical constraints of administering standardized antigen challenges in human studies, direct measurement of parameters such as $K_{1}$, $K_{2}$, $N$. $\mu$, $\eta$, $p_{2}$ and $Q$ is challenging. A viable alternative is to leverage high‑throughput omics profiling combined with large‑scale population datasets to indirectly infer the ensemble averages of these parameters. This approach enables the establishment of stratified baseline references for Immune Capacity $\mathbf{I}$, accounting for influential factors including geography, age, and ethnicity.
6.3 Immune Capacity Assessment in Diseased Populations
In diseased populations, the immune system shifts from maintaining broad potential to targeting specific antigens, engaging the generalized force $Q$ in the response. The extent of immune alteration can be quantified by the work done by $Q$ on the system, defined as:
$I_{\mathrm{Temporal}}=\int_{t_{1}}^{t_{2}}Q(q_{1})\dot{q}_{i}\ \mathrm{d}t$ (31)
Different diseases correspond to distinct disease locations and pathological changes with different immunogenicity $G$, generating different stimulation signals. Consequently, they produce distinct functional forms of $Q$, which ultimately act on the immune system and result in different immune states. Therefore, by comparing immune state differences with the healthy state, the disease course can be determined; and by comparing immune states across different diseases, the disease type can be identified.
6.4 Experimental Measurement of Key Immune Parameters
The theoretical framework establishes that the conserved quantity $I_{\mathrm{Spatial}}$ under homeostasis quantifies the difficulty of altering the immune system’s dynamical state, while $I_{\mathrm{Temporal}}$ quantifies the degree of deviation from homeostasis (i.e., the displacement in $q_{1}$ and the velocity $\dot{q_{1}}$). In practice, comprehensive enumeration of all immune cells is infeasible; therefore, parameters must be estimated from sampled proportions or frequencies. Thus, robust sampling and statistical inference methods are essential. This section outlines potential experimental approaches for measuring these parameters.
6.4.1 Measuring the sensitivity of immune recognition
The quantity $I_{\mathrm{Spatial}}$ governs the sensitivity of the immune response to antigenic challenges and depends on several key parameters: the immunogenicity $G$, the strength of antigenic stimulation $p_{2}$, the total number of immune cells $N$, the per-cell activation barrier $\mu$, the functional and structural coherence of the immune signaling network $\eta$, and the cellular turnover rates $\dot{q_{1}}$.
Experimental estimation of these parameters requires a combination of quantitative assays and mathematical modeling. Immunogenicity $G$ can be assessed through in vitro stimulation assays that measure the dose-response relationship between antigen exposure and immune-cell activation (e.g., proliferation, cytokine secretion). The antigenic impulse $p_{2}$ may be approximated by the magnitude of early immune-cell expansion or antibody titers following controlled antigen exposure. Total immune-cell numbers $N$ are routinely obtained via flow-cytometric immunophenotyping of peripheral blood or lymphoid tissues, coupled with total lymphocyte counts. The activation barrier $\mu$ correlates with the mechanical force threshold for cell triggering, as demonstrated in studies where memory B cells exhibited ultra-low force thresholds due to PIP2-enriched IgG-BCR microdomains[]. The immune-network integrity $\eta$ can be estimated by quantifying the coordination among different immune cell subsets upon stimulation, for example, through the correlation of cytokine secretion profiles across cell types or the synchronicity of activation markers in response to antigen challenge. Finally, the rate of antigenic change $\dot{q_{2}}$ is context-dependent; in acute infections it may be estimated from pathogen replication kinetics (e.g., viral load doubling time), whereas in chronic settings it could correspond to the rate of antigenic drift. Integrating these measurements within a dynamical model allows the composite sensitivity $I_{\mathrm{Spatial}}$ to be estimated, providing a quantitative basis for comparing immune responsiveness across individuals or conditions.
6.4.2 Quantifying Deviation from Homeostasis via Immune Repertoire Sequencing
The extent of immune deviation is measured by contrasting a sample’s receptor repertoire against a reference, which may be another sample or a predetermined homeostatic baseline. Our laboratory previously introduced a computational strategy to define a distance metric between two distinct immune repertoire states. This immunological distance captures both its directional bias (for instance, whether the repertoire shifts toward clonally restricted populations or maintains a balanced, polyclonal composition) and its magnitude (which can be gauged by indicators such as the fraction of germinal-center B cells, the prevalence of activated T cell subsets, and the accrual of somatic hypermutations).
We mathematically framed this distance as the minimum ”cost” required to evolve one immune configuration into another. Conceptually, transforming a state characterized by limited clonality and a low mutational burden into one featuring broad diversity and a higher mutation load entails reshaping the initial clone distribution toward the target profile, while factoring out common elements. This cost is effectively computed using an optimal transport metric, specifically the Wasserstein distance, which quantifies the minimal effort to reconfigure one probability distribution into a different one.
A central consideration is that the biological expenditure for generating clones with varying attributes is not uniform. For example, producing a B-cell lineage carrying three somatic mutations incurs a greater cost than generating one with a single mutation. To account for this, we implemented an in silico, Markov chain-based simulation of repertoire evolution. This model provides empirical estimates for the transition costs between disparate cellular states. The final formulation merges principles from energy landscape theory with optimal transport theory, which portrays the relative ease or difficulty of state transitions and identifies the most efficient evolutionary trajectory. This synthesis bridges abstract computational modeling with tangible immune biology. The framework’s validity was demonstrated through its ability to discriminate mice across varying health strata and to clinically differentiate pediatric febrile illness from Kawasaki disease in a blinded, training-free manner, underscoring its robust predictive capacity and general applicability.
6.5 Sample Size Considerations for Immune Parameter Measurement
When the population size is vastly larger than the sample size, representativeness depends primarily on the absolute sample size needed to overcome stochastic variation and ensure reproducibility, not the sampling ratio. Excessively large sample sizes (e.g., ultra-deep sequencing) entail significant costs with diminishing marginal returns in distributional accuracy. Measurements of adaptive immune system ( $10^{12}$ cells) fall under this premise. Consequently, establishing a population-wide baseline requires a large sample size, but assessing an individual’s status does not.
Conventional flow cytometry (assaying millions of cells) suffices to estimate cell proportions for the adaptive immune system reliably. To stably estimate the distribution shape and compute the difference between distributions of highly diverse immune repertoire ($10^{20}$ unique sequences), which follows a scale-free power-law distribution[], futher mathematical tools are needed, for conventional diversity indices are unsuitable for power-law distributions of immune repertoire[Ref]. Our previous research utilized the Wasserstein distance to compare repertoires, which quantifies the minimal ”cost” to transform one distribution into another. This metric is well-suited for power-law distributions. We demonstrated that approximately $10^{4}$ high-quality sequences are sufficient to stably fit the power-law distribution and enable accurate distance calculation.
7 Microscopic Stochasticity and Macroscopic Immune Patterns
The existence of immunological memory implies that upon returning to homeostasis after an immune response, the potential energy surface of the system to that specific antigen is permanently altered compared to its naive state, manifesting as a reduced average energy barrier against specific antigens. The physical basis of this optimization is somatic hypermutation of antibody genes an inherently stochastic process. The affinity maturation of a B cell clone can be modeled as a random walk within a one-dimensional affinity landscape, where clones with decreased affinity are selected against (cell death), and those with increased affinity are positively selected for expansion. The magnitude of affinity change resulting from each single mutation event follows a Gaussian distribution $N(\mu,\sigma)$ where minute changes are highly probable, while large-affinity leaps or drops are rare. In our model, this stochastic process corresponds to gradually optimizing the antibody structure to refine the system’s evolutionary trajectory, formally adjusting variable $q_{3}$ in system dynamics. Critically, although the specific antibody sequence for any single clone is stochastic, the overall efficiency of the affinity maturation process and the final, macroscopic level of protective immunity achieved are constrained by the system’s key parameters and the conserved quantity $\mathbf{I}$. Thus, idefines the boundary and potential of the system’s statistical behavior.
This inherent stochasticity ensures that even starting from identical initial conditions, the resulting antibody repertoires will be vastly different at the clonal level. This is consistent with experimental and clinical observations that genetically identical twins or inbred mice generate distinct antibody spectra in response to the same pathogen, yet show no significant difference in statistical measures of overall protection, thereby satisfying our model’s symmetry and conservation laws.
The phenomenon of convergent evolution, where similar antibody sequences arise independently against the same epitope (e.g., public clones), provides a bridge from randomness to predictability. This enables the use of artificial intelligence to learn sequence patterns for diagnostic and prognostic applications, although precisely predicting the tertiary structure and binding affinity from the primary amino acid sequence alone remains a formidable challenge.
The same principle of stochastic generation leading to deterministic macroscopic distributions applies to the primary lymphocyte repertoire. Despite utilizing an identical genomic template, stochastic V(D)J recombination generates a highly diverse set of BCR and TCR sequences, the macroscopic statistical properties of which are consistent across individuals.
8 Conclusion and Discussion
In this work, we have established a first-principles theoretical framework for immunity by translating the dynamics of immune recognition into the language of analytical mechanics. By constructing a Lagrangian for the immune system and identifying a continuous symmetry under translation within the antigenic structure space, we have derived a conserved quantity and identified as the Immune Capacity $\mathbf{I}$ via Noether’s theorem. This Immune Capacity $\mathbf{I}$ transforms the phenomenological concept into a rigorously defined, measurable physical invariant, governed by fundamental conservation laws.
The present framework models the immune system as operating in a time‑averaged, steady state and does not explicitly incorporate dynamic regulatory factors such as circadian rhythms or hormonal cycles. These factors are known to modulate immune reactivity on timescales ranging from hours to months and could be formally incorporated as time-varying components of the generalized forces Qacting on the system. By adopting a mean field, steady state representation, we effectively average over such higher frequency or periodic modulations, thereby focusing on the underlying structural and parametric determinants of immune capacity that persist across temporal fluctuations. This simplification is justified when studying coarse-grained, population level differences in immune competence, but future refinements of the model could explicitly introduce oscillatory or episodic forcing terms to capture the dynamic tuning of immune responsiveness by endocrine and neuro immune circuits.
Our theoretical framework is not an abstract construct but is motivated by the need to explain a series of widespread, yet poorly understood, clinical observations. For instance, patients with similar immune status experience comparable disease progression, and that specific antigens (such as certain influenza virus subtypes) can provoke similar symptomatology. These phenomena suggest the existence of a higher-order, conserved principle governing immune responses that transcends specific antigen-receptor interactions. Our work provides a self-consistent, first-principles-based quantitative framework to explain these observations for the first time. Our theoretical framework provides the crucial link to transform these phenomenological observations into testable, quantitative hypotheses.
Our work does not seek to replace existing clinical measures but to explain and unify them under a single theoretical roof. The conserved quantity $I$ provides a physical interpretation for why a ’strong’ or ’weak’ immunity manifests in certain ways. It explains clinical observations as natural consequences of conservation laws such as why a robust system responds calmly to a challenge while a frail system overreact.
The power of this framework lies in its profound unifying and explanatory capacity. Diverse immunological phenomena, from immune memory and tolerance to T cell exhaustion and original antigenic sin, find coherent explanations as specific topological alterations of the underlying potential energy surface (PES), with the system’s evolution obeying the principle of least action. Furthermore, the model reconciles the affinity maturation as a microscopic random walk in sequence space, and link the inherent stochasticity with the deterministic statistical patterns observed at the population level, demonstrating that the conservation of $\mathbf{I}$ sets the boundary conditions for this stochastic evolution.
The quantity $\mathbf{I}$ have two components, $I_{\mathrm{Spatial}}$ and $I_{\mathrm{Temporal}}$, they formally encapsulate the system’s protective sensitivity and response intensity respectively. $(I_{\mathrm{Spatial}}$ possesses the physical dimension of action $\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-1}$, $I_{\mathrm{Temporal}})$ possesses the dimension of energy $\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-2}$.
The implications for applied immunology are significant. Vaccination can be reinterpreted as a strategic engineering of the host’s PES, permanently lowering the energy barrier for a specific pathogen and minimizing the action required for a protective response. The quantity $\mathbf{I}$ thus serves as a system-level metric for vaccine efficacy beyond mere antibody titers. Moreover, the measurement strategies outlined herein pave the way for quantifying immune capacity in clinical settings, moving from qualitative assessments to a physics-based, quantitative score for immune health, with applications in prognostication, immunotherapy monitoring, and gerontology.
Naturally, this initial model has limitations. It treats innate immunity and systemic signals through parameter modulation, and the generalized force $Q$ awaits a more precise mapping onto measurable physiological variables like specific cytokine networks or metabolic fluxes. Future work should focus on refining these connections, expanding the model and ultimately integrating it with other physiological systems such as neuro-endocrine within a whole-organism perspective.
In conclusion, by demonstrating that the immune capacity is a conserved quantity arising from a symmetry principle, this work bridges a fundamental gap between theoretical physics and immunology. It establishes a foundational framework for predictive immunology, where the outcome of host-pathogen encounters and therapeutic interventions can be understood and ultimately predicted from a set of first principles, offering a new perspective of rational immunological design.
Figures
Figures to be completed.
Acknowledgments
Funding
This paper was supported by the National Key R&D Program of China (2023YFC2308200), National Science Foundation for Distinguished Young Scholars (82125037) and the National Natural Science Foundation of China(82230120).
Author contributions:
Yexing Chen: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Project administration, Software, Supervision, Validation, Visualization, Writing - original draft, Writing - review & editing
Qingyun Wei:Formal analysis, Investigation, Validation
Zhongxiang Dong:Investigation
Peng Cao: Conceptualization, Project administration, Investigation, Funding acquisition,Writing-review & editing
Supplementary materials
Supplementary Text
Figs. S
Tables S
Materials and Methods